Circular No. 8662

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 A2 (CATALINA) An apparently asteroidal object discovered by the Catalina Sky Survey

(discovery observation tabulated below), and posted on the ‘NEO Confirmation Page’, has been reported as apparently cometary by several CCD observers, including J. Young (Table Mountain, 0.6-m f/16 Cassegrain reflector; coma diameter ∼ 6′′ in poor seeing conditions on Jan. 23.2 UT), J. E. McGaha (Tucson, AZ, 0.62-m f/5.1 reflector; six stacked 60-s images taken on Jan. 23.2 show a soft 4′′ coma and no tail, but also large changes in brightness of ∼ 0.7 mag over ∼ half an hour), and J. Tichá and M. Tichý (Kleť, 1.06-m KLENOT telescope; diffuse with a coma diameter of 12′′ on images taken on Jan. 23.8).

2006 UT  α_{2000}  δ_{2000}  Mag.
Jan. 21.15860  3^{h}19^{m}54.^{s}37  +25^{o}27′31′′.1  20.3

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-B48.

T = 2005 Aug. 21.780 TT  ω = 152.^{o}246
Ω = 233.244 }2000.0
q = 5.59612 AU  i = 147.894

SUPERNOVAE 2006O AND 2006P Two apparent supernovae have been discovered on unfiltered CCD

frames: 2006O by D. Rich (Hampden, ME, 0.31-m reflector) and 2006P by D. R. Madison and W. Li (LOSS/KAIT; cf. IAUC 8661).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006O Jan. 22.95 1^{h}01^{m}20.^{s}71 +31^{o}30′09′′.4 17.2 6′′.4 W, 18′′.4 S
2006P Jan. 23.22 3 19 17.90 −12 06 18.3 17.5 2′′.9 E, 8′′.8 S

Additional approximate magnitudes of 2006O in UGC 633: 2005 Nov. 19.97 and Dec. 8.99 UT, [18.6; 2006 Jan. 24.09, 17.0 (M. Schwartz). Nothing is visible at the location of 2006O on Palomar Digital Sky Survey images from 1989 (limiting red mag 20.0) and 1988 (limiting blue mag 19.9). Additional KAIT magnitudes of 2006P in MCG –02-9-29: 2005 Dec. 9.21, [19.0; 2006 Jan. 9.22, [18.0; 24.19, 17.4.

2006 January 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green